OpenStudy (anonymous):

The floor of a shed has an area of 88 square feet. The floor is in the shape of a rectangle whose length is 5 feet less than twice the width. Find the length and the width of the floor of the shed. ??????

OpenStudy (anonymous):

how do you find the area of a rectangle?

OpenStudy (anonymous):

l*w

OpenStudy (anonymous):

yes and now how can you say what the length is here using only the width?

OpenStudy (anonymous):

what number is 5 less than twice another number? use W for the other number

OpenStudy (anonymous):

88=2w+(l-5)

OpenStudy (anonymous):

i mean times (l-5)

OpenStudy (anonymous):

close, but we need to get rid of the L... what is L? only using W and numbers and +-*/ ?

OpenStudy (anonymous):

im not sure what your saying

OpenStudy (anonymous):

5 less than twice W

OpenStudy (anonymous):

how can you express that for algrebra?

OpenStudy (anonymous):

2w-5

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

so that is the length and Area = Width * Length so...

OpenStudy (anonymous):

you know the Area, we are using W for the width, and you know the Length so plug them in and solve for W

OpenStudy (anonymous):

ok so 88=W*2W-5

OpenStudy (anonymous):

yes, but be careful here... you need parentheses

OpenStudy (anonymous):

area = (Length)(Width)

OpenStudy (anonymous):

88 = W (2W - 5)

OpenStudy (anonymous):

ok distribute?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

2w^2 - 5w - 88 = 0 solve for W and remember that L = 2W - 5

OpenStudy (anonymous):

ok so then i get -11/2 and 8... well 11*8 equals 88 but the

OpenStudy (anonymous):

then it doesnt match and idk im lost

OpenStudy (anonymous):

forget the negative answer... go with the 8

OpenStudy (anonymous):

the width is 8 so the length is 2(8)-5